Staind, Inc., has 8 percent coupon bonds on the market that have 14 years left to maturity. The bonds make annual payments. If the YTM on these bonds is 9 percent, what is the current bond price?


rev: 09_18_2012

$1,030.00

$922.14

$1,086.44

$876.03

$968.25
Answer #1

Annual coupon=1000*8%=80

Hence current price=Annual coupon*Present value of annuity factor(9%,14)+$1000*Present value of discounting factor(9%,14)

=80*7.786150389+$1000*0.299246465

=$922.14(Approx).

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=80[1-(1.09)^-14]/0.09

=80*7.786150389

2.Present value of discounting factor=1000/1.09^14

=1000*0.299246465